if a[1]=4 and d=-2, what are the first four terms of the arithmetic sequence?
Answer by stanbon(75887)

if a[1]=4 and d=-2, what are the first four terms of the arithmetic sequence?
4,4-2=2, 2-2=0, 0-2=-2
==================================
Cheers,
Stan H.